OpenStudy (anonymous):

Tokyo-Moto Quad Racers found that its daily profit, P, in thousands, is given by P = -8x2 + 96x - 256 where x is the number of Quad Racers manufactured daily. Which of the following is the best estimate for the maximum profit the manufacturer can earn per day, in thousands?

OpenStudy (anonymous):

find the vertex of the parabola. first coordinate of the vertex is given by \(-\frac{b}{2a}\)

OpenStudy (anonymous):

answer choices are A. 32 B. 35 C. 40 D. 38

OpenStudy (anonymous):

your job is to compute \(-\frac{b}{2a}\) for \(a=-8,b=96\)

OpenStudy (anonymous):

@satellite73 so out of the answers choices i listed which would be the one

OpenStudy (anonymous):

when you do that, you will get 6 replace x by 6 and find the corresponding P

OpenStudy (anonymous):

\[-\frac{96}{2\times (-8)}=\frac{96}{16}=6\] \[P=-8\times 6^2+96\times 6-256\]

OpenStudy (anonymous):

i get 32 but you should check my arithmetic

OpenStudy (anonymous):

@satellite73 thank you i love you!

OpenStudy (anonymous):

An architect's model of a building uses a scale in which 1/2 inch equals 1 foot. If the width of the building is 60 feet, how many inches will the width be on the model?